Solve 2x2 - x - 6 = 0 by completing the squares.

Answer by Alan3354(69443)

Solve 2x2 - x - 6 = 0 by completing the squares.
--------
2x2 - x - 6 = 0
Use ^ (Shift 6) for exponents.
---
2x^2 - x - 6 = 0
---
x^2 - x/2 = 3
Find c
c = (-b/2a)^2 = ((-1/2)/2)^2 = 1/16
Add 1/16 to each side.
x^2 - x/2 + 1/16 = 3.0625
(x - 1/4)^2 = 1.75^2
x - 1/4 = 1.75
x = 2
---------------
x - 1/4 = -1.75
x = -1.5